### v1

# Merging of Kolmogorov-Arnold networks trained on disjoint datasets

Abstract
Training on disjoint datasets can serve two primary goals: accelerating data processing and enabling federated learning. It has already been established that Kolmogorov-Arnold networks (KANs) are particularly well suited for federated learning and can be merged through simple parameter averaging. While the federated learning literature has mostly focused on achieving training convergence across distributed nodes, the present paper specifically targets acceleration of the training, which depends critically on the choice of an optimisation method and the type of the basis functions. To the best knowledge of the authors, the fastest currently-available combination is the Newton-Kaczmarz method and the piecewise-linear basis functions. Here, it is shown that training on disjoint datasets (or disjoint subsets of the training dataset) can further improve the performance. Experimental comparisons are provided, and all corresponding codes are publicly available.
Keywords: Newton-Kaczmarz method, piecewise-linear functions, training on disjoint datasets.
1  Introduction
Since the publication of Liu et al. [1], which advocated spline-based univariate functions together with the ‘Adam’ or LBFGS optimisers for the training, these design choices have become dominant in the recent literature on Kolmogorov-Arnold networks (KANs). Several works have proposed alternative basis functions, aimed at reducing the computational overhead relative to neural networks, such as radial or alternative polynomial bases. However, these approaches remain conceptually similar — they rely on smooth basis functions [^1] Term ‘smooth’ is used loosely, implying the existence of some number of continuous derivatives; formally, the term of course means belonging to $C^{\infty}$ class. and typically require a relatively large number of calculations per each underlying function of KAN. Earlier alternatives, predating [1], include ReLU-type and piecewise-linear basis functions [2, 3], which were observed to be faster, but most subsequent research followed [1] as the principal reference point. It is important to emphasise that KANs were not invented in 2024 and originate from the pioneering work of Igelnik and Parikh [4] published in 2003. A historical overview of KANs’ development can be found in e.g. [5] and references therein.
A number of recent studies have evaluated the federated learning performance of KANs relative to multilayer perceptrons (MLPs) [6, 7, 8, 9]. These works also adopt the spline-based architectures and various optimisation methods. While improved accuracy for KANs is consistently reported, the computational cost is significantly higher. For example, the authors of [6] report training times of $314\,\text{s}$ for their KAN variant vs. $2.35\,\text{s}$ for an MLP on the same task. In the present paper, it is shown that such difference is not intrinsic to the KAN model class; rather, it arises from particular choices of basis functions and optimisation algorithms. For the piecewise-linear basis functions and the Newton-Kaczmarz optimisation scheme, this disparity is reversed, and highly-efficient KANs’ training is achieved while maintaining comparable accuracy. Moreover, in the present paper, for the first time, it is shown that the application of the federated-learning idea, i.e. training on disjoint subsets of a training dataset and merging the models, can significantly accelerate the training and represents a form of parallelisation of the training.
Although basis functions with continuous derivatives are essential for some applications, such as physics-informed modelling, they are not universally required. Moreover, the basis used during the training does not need to coincide with the basis used in the final model. In particular, a model may be trained efficiently using a piecewise-linear representation and subsequently upgraded to splines or other smooth bases with only minor post-training, thereby preserving accuracy while substantially reducing the training time. Thus, selecting piecewise-linear functions does not represent a move towards a more primitive model; instead, it enables faster optimisation while retaining full descriptive capabilities. The latter is further elaborated in [5].
2  Approach
A Kaczmarz-based training method for KANs with piecewise-linear basis functions was introduced by the authors in 2020 in a preprint later published as [3]; the approach was generalised to the Newton-Kaczmarz training method for KANs with arbitrary basis functions in 2023 in a preprint later published as [5]. The reader is referred to the above publications for details; however, for completeness of the paper, a brief summary is provided below, assuming that the reader is familiar with KANs.
The basic building block of KAN is an operator that converts $n$-dimensional vector $\boldsymbol{y}$ into scalar $\hat{z}$. It can be called the discrete Urysohn operator or a generalised additive model:
$\hat{z}=\sum_{j=1}^{n}g_{j}\left(y_{j}\right),$ (1)
where $g_{j}$ are some functions. A layer of KAN consists of a set of such blocks, forming $m$-dimensional vector $\hat{\boldsymbol{z}}$, such that
$\hat{z}_{i}=\sum_{j=1}^{n}g_{i,j}\left(y_{j}\right),\quad i=1,2,\ldots,m.$ (2)
Thus, a layer converts vector $\boldsymbol{y}$ into vector $\hat{\boldsymbol{z}}$. KAN is a chain of such layers, where the outputs of the previous layer are the inputs for the next layer.
Here, functions $g_{i,j}$ are taken to be piecewise-linear. Each argument $y_{j}$ lies within a single linear segment. For uniformly-spaced nodes with inter-nodal distance $\Delta y$, the interpolation is trivially computed via
$s=1+\left(y_{j}-y_{\mathrm{min}}\right)/\Delta y,\qquad k=\lfloor s\rfloor,\qquad f=s-k,$ (3)
$g_{i,j}\left(y_{j}\right)=\left(1-f\right)G_{i,j,k}+fG_{i,j,k+1},$ (4)
where $k$ is the index of the left node, $f$ is the relative offset from that node, and $G_{i,j,k}$ are the model’s parameters (values of $g_{i,j}$ at the nodes).
Training of KAN consists in updating parameters layer-by-layer for each record of the training dataset. To update one layer, a ‘desired’ (i.e. ‘target’) vector $\boldsymbol{z}$ is required. One can execute such update by performing one ‘damped’ step of a gradient descend, applied to layer loss function $L=\left\lVert\boldsymbol{z}-\boldsymbol{\hat{z}}\right\rVert^{2}$, resulting in
$G_{i,j,k}^{\mathrm{new}}=G_{i,j,k}^{\mathrm{old}}+\mu\left(z_{i}-\hat{z}_{i}\right)\left(1-f\right),$ (5)
$G_{i,j,k+1}^{\mathrm{new}}=G_{i,j,k+1}^{\mathrm{old}}+\mu\left(z_{i}-\hat{z}_{i}\right)f,$ (6)
where $k$ and $f$ depend on index $j$ according to equation (3), multiplier $\mu$ is the scaling factor, which depends on all layers in the original form of the method.
For the outer layer of KAN, the target vector is given in the training dataset. For some given layer, if target vector $\boldsymbol{z}$ is known, then one can calculate target vector $\boldsymbol{y}$ for the preceding layer simply by the chain rule for the derivatives:
$\Delta\boldsymbol{y}=J^{\mathrm{T}}\Delta\boldsymbol{z},$ (7)
where $\Delta\boldsymbol{z}=\boldsymbol{z}-\boldsymbol{\hat{z}}$ and $\Delta\boldsymbol{y}=\boldsymbol{y}-\boldsymbol{\hat{y}}$ are the residuals; matrix $J$ consists of partial derivatives, such that $J_{i,j}=\partial g_{i,j}/\partial y_{j}$, which are trivial to compute for piecewise-linear functions.
3  Merging of the models
The described approach iterates record-by-record through the training dataset. The hypothesis of the present paper is that such training can be parallelised using the algorithm as follows.
A number of equal-size subsets (batches) are formed using the records from the training dataset. Each batch should be sufficiently representative of the modelled system, i.e. it is implied that the algorithm will be used for large data. Hence, one should choose the number of batches and the batch size.
For each batch, an exact copy of the model (KAN) is created and trained on the corresponding batch (one pass through the data). The models are trained in parallel. After the training is finished, the models are merged into a single model simply by computing the mean value of every parameter.
The above steps (forming batches, creating copies of the model, training copies, merging copies) are iteratively repeated until the selected convergence criteria is reached. The performance of this approach should of course depend on the number of batches and on the batch size. In contrast to federated learning, these parameters may be chosen arbitrarily.
Using this approach, processing of the records from the training dataset should be parallelisable almost ideally, with minor losses for model coping/merging. However, there will be some loss of accuracy due to the merging, and one may expect increasing error with the number of batches. Therefore, there should be an optimal number of batches. Thus, the computational experiments serve two completely separate purposes: the first is to establish the scaling of the implementation of the algorithm (i.e. how much time is lost for coping/merging), the second is to find how the accuracy drops with the number of batches.
The loss of accuracy due to the merging is related to the parallel models all being equal in the beginning of the training on the batches. However, in the training algorithm, the model’s update depends on the current model’s parameters. Thus, in the sequential case, each update is performed for the up-to-date model, while in the parallel case, threads starting from the second work with the ‘old’ model.
4  Experimental part performed on laptop — strong scaling
4.1  Baseline
The single-threaded implementation of the approach described in section 2 has already been extensively tested in the previous publication by the authors. It has been shown that it significantly outperforms state-of-the-art MATLAB’s neural networks. However, for completeness of the present paper, it is worth re-establishing a baseline result — showing a brief comparison between the considered single-threaded (sequential) implementation and MATLAB’s neural networks.
Tests are conducted on a CPU (Intel(R) Core(TM) Ultra 7 155H, 16 cores) and on a GPU (Nvidia GeForce RTX 4050 Laptop). Obviously, one cannot objectively compare a CPU code with a GPU code, as they run on completely different hardware. However, the authors believe that their setup is representative of a typical modern researcher’s laptop, and it is beneficial to see the CPU/GPU difference in this particular case, i.e. whether performance gain of MATLAB’s neural networks when switching to a GPU is sufficient to overtake the authors KANs’ implementation on such particular hardware.
For the tests, synthetic data is generated. The inputs are elements of $4\times 4$ matrices and the outputs are the determinants of the matrices. The training dataset contains $100\mathrm{K}$ records; the validation dataset contains $20\mathrm{K}$ records. Random data is generated each execution of the code. The accuracy metric is the Pearson correlation coefficient (in percent) between the real and the modelled output.
This is just a demonstrative example. Therefore, only one configuration of MATLAB’s neural network is used, with $5$ hidden layers, each consisting of $10$ neurons (i.e. the total of $621$ parameters per model); only one KAN configuration in the authors’ code is used, with $70$ addends (classical two-layer architecture), $3$ points per inner function, and $25$ points per outer function (i.e. the total of $5,110$ parameters per model). Readers who are interested in a more detailed comparison with varying MLP and KAN architectures are referred to the previous paper [5].
The results are shown in table 1, where it is seen that MLPs are outperformed by the authors’ implementation of KANs. When switching to a GPU, MATLAB’s neural networks are approximately $3$ times faster (on the particular tested hardware), but still have slightly lower accuracy and longer training time than KANs.
MLP, CPU MLP, GPU KAN
Pearson (%) $94.6\pm 1.4$ $95.0\pm 0.9$ $97.1\pm 0.1$
time (s) $32.7\pm 0.9$ $8.6\pm 0.3$ $5.9\pm 0.6$
Table 1: Baseline accuracy and performance tests for modelling of the determinants of random $4\times 4$ matrices conducted for MATLAB’s neural networks and the authors’ KANs’ implementation (single-threaded, C++ implementation for CPU). The values are expectations and standard deviations for 88 executions.
4.2  Scaling test
In this section, the main hypothesis of the paper that parallelisation can be implemented by training on disjoint batches across multiple threads is benchmarked. The strong scaling is tested: the total computational work is fixed, which is the total number of processed records of the training dataset, while the work is equally split between the threads. The batch size and the number of threads are varied. As usual, the speedup is measured as $S=t(1)/t(n)$, where $t(1)$ is the time required to complete the work on $1$ thread and $t(n)$ is the time required to complete the same work on $n$ threads.
In the single-threaded setting for predicting determinants of $4\times 4$ matrices, approximately $10$ passes through the training dataset (epochs) are required to reach the Pearson correlation of at least $97\%$. This corresponds to processing $10^{6}$ records, and this value is used as the fixed workload for all parallel configurations. The number of threads is varied from $1$ to $6$ (the particular Intel CPU has $6$ performance ‘fast’ cores and $10$ efficiency ‘slow’ cores). Batch sizes range from $50\mathrm{K}$ down to $3\mathrm{K}$ records. For each combination of thread count and batch size, the required number of model merges (rounds) is chosen such that the total number of processed records remains exactly $1\mathrm{M}$.
The results are shown in table 2, where it is seen that the scaling is sufficiently close to linear, given that the experiments are performed on a Windows laptop (i.e. there is little control over background activity, and the hardware is not as optimised for parallel execution as the HPC hardware). The speedup on $4$ threads is between $3.7$ and $4.1$, depending on the batch size, while the speedup on $6$ threads is between $4.5$ and $4.9$.
Under ideal scaling, the accuracy would remain unchanged; however, the merging of the models leads to a computational error, and the actual value drops from $96.8\%$ for $1$ thread to $94.5\%$ for $6$ threads. Such loss in accuracy can be compensated of course by increasing the number of rounds. For example, using $6$ threads, $20$ rounds, and the batch size of $16,666$ (twice larger computational work than in the table) yields the Pearson correlation of $97.0\%$, while completing in $2.22$ seconds — $2.4$ times faster than for $1$ thread.
threads rounds batch size Pearson (%) time (s)
1 10 100,000 96.8 5.4
2 10 50,000 96.7 2.58
2 20 25,000 96.6 2.44
2 40 12,500 96.3 2.63
3 10 33,333 96.6 1.83
3 20 16,666 96.0 1.86
3 40 16,666 96.1 1.79
4 10 25,000 96.1 1.32
4 20 12,500 95.7 1.42
4 40 6,250 95.8 1.46
5 10 20,000 95.5 1.21
5 20 10,000 94.9 1.27
5 40 5,000 95.1 1.20
6 10 16,666 94.5 1.10
6 20 8,333 94.4 1.12
6 40 4,166 94.6 1.19
Table 2: Accuracies and execution times of the code on a laptop for the example of matrix determinants — strong scaling test.
4.3  Second example
The second example is not strictly a scaling test, as the computational workload varies, but it serves as a benchmark of parallel training to a fixed accuracy. The training is done on synthetic data, where the outputs are the areas of triangular faces of randomly-generated tetrahedra ($4$ faces), and the inputs are the coordinates of the vertices of the tetrahedra in 3D ($4$ vertices, $3$ coordinates for each). Number of the training records is $500\mathrm{K}$, number of the validation records is $50\mathrm{K}$. Three-layered KANs are trained, the first layer consisting of $70$ operators with $3$ points per function, the second layer consisting of $30$ operators with $18$ points per function, the third layer consisting of $4$ operators with $22$ points per function (i.e. the total of $42,960$ model parameters). MATLAB’s neural networks are used for the comparison with $4$ hidden layers and $80$ neurons per layer ($20,804$ model parameters). The results are shown in table 3, where it is seen that when trained to the same accuracy, parallel implementation of KANs is about $2.8$ times faster than the sequential and two orders of magnitude faster than MLP training on CPU.
Pearson (%) for outputs time (s)
MLP, CPU $98.6\pm 0.7$ $98.2\pm 0.9$ $98.9\pm 0.3$ $97.9\pm 0.4$ $1126.4\pm 69.5$
MLP, GPU $98.7\pm 0.5$ $98.3\pm 0.7$ $98.2\pm 0.8$ $98.3\pm 0.9$ $76.9\pm 1.2$
KAN sequential $8.6\pm 0.6 99$$9.0\pm 0.4$ $98.1\pm 0.2$ $98.5\pm 0.3$ $29.2\pm 3.2$
KAN parallel, 6 threads $98.5\pm 0.9$ $98.1\pm 0.7$ $98.6\pm 0.3$ $98.3\pm 0.4$ $10.5\pm 1.3$
Table 3: Baseline accuracy and performance tests for modelling of the areas of faces of random tetrahedrons conducted for MATLAB’s neural networks and the authors’ KANs’ implementation.
5  Experimental part performed on cluster — weak scaling
To assess the scalability of the code, the tests were run on an HPC cluster. All jobs were run on a single node equipped with two Intel(R) Xeon(R) Platinum 8358 CPUs ($32$ cores per CPU) and $256\mathrm{GB}$ RAM. The tests consisted in training KANs on datasets with the inputs being elements of $4\times 4$ or $5\times 5$ matrices and the outputs being the determinants of the matrices.
In this section, the weak scaling is tested: the number of threads was kept proportional to the total number of passes through the data (epochs), i.e. the total computational work. Obviously, in the case of ideal (linear) scaling, the execution times should be the same. As usual, the efficiency is measured as $E=t(1)/t(n)$, where $t(1)$ is the time required to complete work on $1$ thread and $t(n)$ is the time require to complete work that is $n$ times larger on $n$ threads.
The results are shown in table 4, where it is seen that the scaling is very close to linear. For the example of $4\times 4$ matrix determinants, the efficiency is above $93\%$ when up to $32$ threads are used. For the example of $5\times 5$ matrix determinants, the efficiency is above $95\%$ when up to $16$ threads are used and only drops down to $83\%$ when $32$ threads are used (and even further for $64$ threads). Such noticeable decline in the latter case is due to a large number of model synchronisations and large model sizes for $5\times 5$ example.
In the example of $4\times 4$ matrix determinants, the datasets consisted of $100\mathrm{K}$ training records, the models consisted of $70$ addends, with $3$ points per inner function and $25$ points per outer function (i.e. the total of $5,110$ parameters per model). The batch size was taken to be $100\mathrm{K}$ records, i.e. the entire training dataset. The jobs consisted of $10$ batches, i.e. $10$ synchronisations (of $n$ models, where $n$ is the number of threads) are made within each execution of the code.
In the example of $5\times 5$ matrix determinants, the datasets consisted of $10\mathrm{M}$ training records, the models consisted of $160$ addends, with $3$ points per inner functions and $30$ points per outer function (i.e. the total of $16,800$ parameters per model). The batch size was taken to be $50\mathrm{K}$ records. The jobs consisted of $200$ batches.
threads 11 22 44 88 $16$ $32$ $64$
time (s), $4\times 4$ $6.64$ $6.80$ $6.82$ $6.82$ $7.11$ $7.04$ $7.79$
time (s), $5\times 5$ $217.02$ $220.24$ $226.07$ $222.82$ $227.68$ $260.99$ $282.96$
Table 4: Execution times of the code on HPC cluster for examples of matrix determinants — weak scaling test. The values are averages of three programme runs.
6  Concurrent-serial alternations of Newton-Kaczmarz
The conceptual description of the Newton-Kaczmarz algorithm provided in section 2 highlights its alternating concurrent-sequential structure. Each layer maps an input vector to an output vector through multiple independent operations that may be executed concurrently. However, the outputs of all concurrent computations must be synchronised before proceeding to the next layer. The backward updates follow the same pattern — a chain of potentially concurrent steps separated by unavoidable sequential dependencies.
These serial steps between independent computations represent a significant obstacle for concurrent training. Neither CPU nor GPU kernels are well suited for such patterns. This limitation can be illustrated using the classical vector-input-scalar-output two-layer KAN:
$\hat{z}=\sum_{j=1}^{n}A_{j},\qquad A_{j}=g_{j}\left(\sum_{l=1}^{p}f_{j,l}\left(x_{l}\right)\right),$ (8)
where $A_{j}$ stands for one addend, and each addend is itself a small tree mapping inputs $x_{l}$ to a scalar. All $n$ addends must be computed in order to obtain the residual $\Delta z$, and these computations can be executed concurrently. This residual, divided evenly among the addends and scaled by the learning rate, is the only quantity required for updating all functions $g_{j}$ and $f_{j,l}$. In the piecewise linear model, evaluating a function requires only two operations (one division and one multiplication), which is far too little computation for concurrency on CPUs or GPUs to provide any advantage. The synchronisation required after all $n$ addends are computed often takes longer than the time saved through parallelism.
However, this pattern is well suited for customised hardware, such as field-programmable gate arrays (FPGAs) or application-specific integrated circuits (ASICs). A CPU-based simulation demonstrating a custom synchronisation barrier designed by the authors is available online [^2] https://github.com/andrewpolar/detfast, but it serves only as a conceptual demonstration; synchronising at each training record introduces substantial overhead. The example implemented in the code is training a model for determinates of $5\times 5$ matrices. Training MLPs for the latter requires several hours [5], whereas the CPU-based KAN code completes it in a few minutes. An FPGA or ASIC implementation could reduce this further to seconds, because each addend can be mapped to a dedicated hardware processing block. In such customised hardware, concurrency is not emulated through threads or kernels but is realised through physical parallelism, where synchronization and data flow are governed directly by wiring rather than by software-controlled barriers.
6.1  Pretraining
Structure (8) motivates a pretraining strategy, in which groups of addends are trained concurrently over the entire dataset to predict an equal share of target $z$. For the example of determinants of $5\times 5$ matrices and a dataset of $10\mathrm{M}$ records, such pretraining produces a model with the Pearson correlation near $70\%$, providing a strong initialisation for the main training phase. When pretraining is combined with the concurrent-serial alternations, the full training process for $5\times 5$ determinants completes in just over $2$ minutes on the CPU used for the baseline examples.
7  Conclusions
The results demonstrate that KANs’ training can indeed be accelerated by iteratively training copies of the model on disjoint datasets (batches) and merging them. The speedup ratios between sequential and concurrent KAN training typically fall in the range of 3344, meaning that acceleration is possible, but there are some limits.
At the same time, even the sequential KAN implementation by the authors consistently outperforms MLPs on CPUs and, remarkably, also on a GPU. While direct CPU-GPU comparisons are formally inappropriate due to differences in hardware, these results are included for reference, taking into account that the GPU used for the benchmarking can be considered to be a typical laptop GPU (it has been noticed by the authors that some researchers have unrealistic expectations from GPU implementations).
Since methodology adopted in the authors’ code is FPGA-friendly, the authors believe that this direction is a perspective research capable to reduce hours into seconds of processing time for very large models and datasets.
MATLAB scripts [^3] http://openkan.org/downloadskan/matlabscripts.zip, sequential implementation [^4] https://github.com/andrewpolar/kankan-8, and parallel implementation [^5] https://github.com/andrewpolar/meanmodel are available online.